# TriDi: Trilateral Diffusion of 3D Humans, Objects, and Interactions

### v2

@@ 9 Additional Evaluation
This section introduces details on the user study that was used to evaluate TriDi. We have designed and run a user study, asking participants to rate the quality of the generated interactions. We compared TriDi against one baseline method and ground-truth data in two generation modes: $p({\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{H}}},{\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{I}}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}{\mathcal{O}}})$ and $p({\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{O}}},{\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{I}}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}{\mathcal{H}}})$. We used GNet and ObjPOP+cVAE as the baselines, and randomly selected 10 queries for the generation (5 from each of BEHAVE and GRAB) for each mode. In every question we show users three randomly shuffled samples: ground-truth data, TriDi, and corresponding baseline. The participants were asked to rate the quality of each sample based on the realism of human-object interaction, and the amount of interpenetration between human and object. Each sample is rendered from the same 4 orthogonal views to allow comprehensive assessment . The rating scale consisted of three options: Worst, Moderate, and Best, with ratings being non-exclusive (i.e., more than one sample can have a similar rating). Example interface of the user study is provided in the Fig. S3. As a result, we have collected 40 responses. We summarize the results in the Tab. S2, comparing the ratings assigned to the samples by users. On average, results of TriDi were preferred to the baselines in $89.0\%$ of the cases and preferred to the ground-truth examples in $52.0\%$ of the cases. This suggests that the results of TriDi are more appreciable than the baselines and produce a realism comparable to captured data.
@@
$p({\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{H}}},{\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{I}}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}{\mathcal{O}}})$ TriDi $>$ GNet $87.75\%$
TriDi $>$ GT data $47.75\%$
$p({\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{O}}},{\color[rgb]{0.0625,0.4765625,0.07421875}{\mathcal{I}}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}{\mathcal{H}}})$ TriDi $>$ ObjPOP+cVAE $90.25\%$
TriDi $>$ GT data $56.25\%$
@@
NoGuide $\underline{21.5}$ $\underline{12.4}$ $\mathbf{96.0}/\mathbf{96.5}$ $\underline{28.1}$ $\underline{15.4}$ 96.2/$\mathbf{96.2}/\mathbf{96.2}$
$(\mathcal{H},\mathcal{O})$ 21.9 12.7 $\mathbf{96.0}$ / NA $28.4$ 15.6 $\underline{96.1}$ / NA
NoAug 23.2 12.9 95.4 / $\underline{96.2}$ 31.0 17.8 95.5 / $96.0$
@@
NoGuide $16.2$ $11.3$ $97.5$ / $\underline{98.3}$ $9.0$ $7.5$ $98.2$ / $\underline{98.3}$
$(\mathcal{H},\mathcal{O})$ $17.3$ $11.8$ $97.3$ / NA $9.5$ $7.9$ $98.0$ / NA
@@ 11 Datasets
BEHAVE [3] captures 8 subjects interacting with 20 different objects, represented as SMPL+H meshes and global configuration, respectively. We downsample the $30fps$ train sequences to $10fps$ and consider the official $1fps$ test subset.
@@
We use the subset of GRAB [67] introduced in [59]. This subset includes 10 subjects interacting with 20 objects. The $120fps$ train and test sequences are downsampled to $1fps$. The test set consists of interactions performed by subjects 9 and 10.
@@
We downsample the original $30fps$ sequences to $10fps$ and follow the train-test split provided by VisTracker [81]: Data from subjects 1-8 is used for training, and sequences from subjects 9 and 10 are used for evaluation.
@@ 13 Error Metrics
Diversity (Div) [22]. Diversity measures the variance of the generated samples. Two subsets $S_{1}=\{v_{1},...,v_{|S|}\}$ and $S_{2}=\{v^{\prime}_{1},...,v^{\prime}_{|S|}\}$ of the same size $|S|=200$ are drawn from either $S_{g}$ or $S_{r}$ (depending on whether we want to evaluate the metric for the method or the GT data). The diversity then is computed as follows:
@@
Multimodality (MMod) [22]. Multimodality measures the variance of the generated samples within the same object category. For every object class $c\in{1,...C}$ two subsets $S^{c}_{1}=\{v_{c,1},...,v_{c,|S|}\}$ and $S^{c}_{2}=\{v^{\prime}_{c,1},...,v^{\prime}_{c,|S|}\}$ of the same size $|S|=200$ are drawn from either $S_{g}$ or $S_{r}$. The multimodality is then computed as follows ($S_{1}=\{S^{1}_{1},...,S^{C}_{1}\},S_{2}=\{S^{1}_{2},...,S^{C}_{2}\}$):

### v3

@@ 9 Additional Evaluation
[…]
@@
[…]
@@
Evaluation of ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}},{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{I}}$.
We compare TriDi with COINS [109] on the task of human generation given object and text in Table S3. We observe that while COINS is able to generate sitting poses it struggles to generate realistic and diverse interactions with other objects.
Method BEHAVE, ${\color[rgb]{0.0625,0.4765625,0.07421875}\mathcal{H}}|{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{O}},{\color[rgb]{0.05078125,0.40234375,0.59765625}\mathcal{I}}$
[…]
COINS $96.4^{\pm 0.1}$ $19.6^{\pm 0.1}$ $3.02^{\pm 0.008}$ 43.4 15.9 $\nicefrac{{93.9}}{{\textnormal{NA}}}$
TriDi $\mathbf{66.1^{\pm 0.4}}$ $\mathbf{50.8^{\pm 0.1}}$ $\mathbf{1.30^{\pm 0.010}}$ $\mathbf{16.9}$ $\mathbf{10.6}$ $\nicefrac{{\mathbf{96.7}}}{{\mathbf{99.5}}}$
Table S3: Comparison with COINS.
@@
NoGuide $\underline{21.5}$ $\underline{12.4}$ $\mathbf{96.0}/\mathbf{96.5}$ $\underline{28.1}$ $\underline{15.4}$ $\mathbf{96.2}/\mathbf{96.2}$
$(\mathcal{H},\mathcal{O})$ 21.9 12.7 $\mathbf{96.0}$ / NA 28.428.428.4 15.6 $\underline{96.1}$ / NA
NoAug 23.2 12.9 95.4 / $\underline{96.2}$ 31.0 17.8 95.5 / 96.096.096.0
@@
NoGuide 16.216.216.2 11.311.311.3 97.597.597.5 / $\underline{98.3}$ 9.09.09.0 7.57.57.5 98.298.298.2 / $\underline{98.3}$
$(\mathcal{H},\mathcal{O})$ 17.317.317.3 11.811.811.8 97.397.397.3 / NA 9.59.59.5 7.97.97.9 98.098.098.0 / NA
@@
Evaluation of penetration.
We compute SDF-based penetration metrics in Tab.S8: average min. dist. between H and O (Min. D. $[cm.]$), contact percentage and penetration score ($C_{\%},P_{sc}[cm.])$, CHOIS [109]), penetration depth ($PD[cm.]$, DiffH2[14]). TriDi’s results are close to values computed for ground-truth data, outperforming the baselines that generate floaters and more penetrations.
BEHAVE
Min. D.$\downarrow$ $C_{\%}\uparrow$ $PD\downarrow$ $P_{sc}\downarrow$ Min. D.$\downarrow$ $C_{\%}\uparrow$ $PD\downarrow$ $P_{sc}\downarrow$
Data $\mathbf{0.65}$ $\mathbf{99.6}$ $\mathbf{1.73}$ $\mathbf{0.03}$ $\underline{0.65}$ $\mathbf{99.6}$ $\mathbf{1.73}$ $\mathbf{0.03}$
ObjPOP + cVAE - - - - 1.25 92.8 2.8 0.07
GNet 1.50 91.9 3.6 0.08 - - -
TriDi (Ours) $\underline{0.74}$ $\underline{96.5}$ $\underline{2.6}$ $\underline{0.07}$ $\mathbf{0.56}$ $\underline{97.7}$ $\underline{2.38}$ $\underline{0.05}$
GRAB
Min. D.$\downarrow$ $C_{\%}\uparrow$ $PD\downarrow$ $P_{sc}\downarrow$ Min. D.$\downarrow$ $C_{\%}\uparrow$ $PD\downarrow$ $P_{sc}\downarrow$
Data $\mathbf{0.08}$ $\mathbf{100.0}$ $\underline{0.46}$ $\underline{0.0012}$ $\mathbf{0.08}$ $\mathbf{100.0}$ $\underline{0.46}$ $\underline{0.0012}$
ObjPOP + cVAE - - - - 4.98 60.1 $\mathbf{0.17}$ $\mathbf{0.0005}$
GNet 5.99 58.9 $\mathbf{0.18}$ $\mathbf{0.0002}$
TriDi (Ours) $\underline{0.26}$ $\underline{99.3}$ 0.79 0.0052 $\underline{0.29}$ $\underline{98.8}$ 0.81 0.0048
Table S8: Penetration analysis.
@@ 11 Datasets
[…]
@@
[…]
@@
[…]
@@ 13 Error Metrics
[…]
@@
[…]